OpenStudy (frostbite):

Just need a minor reminder about how we write the domain of a function with good old mathematical notation.

OpenStudy (anonymous):

it is really how you write an interval all real numbers would be either \(\mathbb{R}\) or \((-\infty, \infty)\)

OpenStudy (anonymous):

all number greater than or equal to -1 (the domain of \(f(x)=\sqrt{x+1}\) ) would be either \[x\geq-1\] or \([-1,\infty)\)

OpenStudy (anonymous):

if you want to be fancy and use set builder notation, you could write \[\{x:x\geq -1\}\]

OpenStudy (frostbite):

Was more thinking something like this (i know this might just be wrong but still): Df={x,y│x,y∈4xy-3y^2>0}

OpenStudy (anonymous):

What's that, x and y a function of z?

OpenStudy (frostbite):

Okay i just give the full task: A function f is given by f(x,y)=sqrt(4xy-3y^2). I am going to find the domain, but i wanted to start writeing the domain with mathematical notation and a bit unsure about how it look lie if it is Df={f(x)│x,y∈4xy-3y^2>0} or how it was again

OpenStudy (frostbite):

look like*

OpenStudy (frostbite):

Df is just the domain in my language

OpenStudy (anonymous):

{x,y│4xy-3y^2>0} (greater than or equal to 0 is also OK)

OpenStudy (anonymous):

x,y in R?

OpenStudy (frostbite):

Thanks what i was looking for :)

OpenStudy (anonymous):

{x,y in R│4xy-3y^2>0}

OpenStudy (anonymous):

Where I come from ,this would be understood without specification.....

OpenStudy (frostbite):

yea or we can be very annoying and write {∀x,y∈C}?

OpenStudy (anonymous):

x,y in C is sufficient

OpenStudy (anonymous):

Let me just think, complex square root.....

OpenStudy (anonymous):

Yes, principal value...

OpenStudy (frostbite):

Ok but can we then write the following: {x,y∈R│4xy-3y^2>0} i mean the ∈ just means "belong to" right?

OpenStudy (anonymous):

Yes, or simply "in"

OpenStudy (frostbite):

Well i say thanks. Might write a new question when i need to solve the inequality.

OpenStudy (anonymous):

ur welcome.